A square loop 10 cm on a side consists of 100 turns of wire that experiences a minimum torque of zero and a maximum torque of 0.045 Nm. If the current in the loop is 2.82 A, calculate the magnetic field strength that the loop rests in
Answer #1

Area

A=S^2=0.1^2=0.01 m^2

Maximum torque

Tmax=NIAB

=>B=Tmax/NIA

B=0.045/100*2.82*0.01

B=0.01596 T or 15.96 mT